OpenStudy (anonymous):

Solve the formula for the given letter. Assume all variable represent nonnegative numbers. a^2+n^2=h^2, for a

OpenStudy (anonymous):

\[a^2+n^2=h^2\] \[a^2=h^2-n^2\] \[a=\sqrt{a^2-n^2}\]what myonlybro said

OpenStudy (anonymous):

? it disappeared!

OpenStudy (anonymous):

what disappeared?